Mark and Brian are starting a new business. They borrowed $7,500 from their bank in order to help with start-up costs. If the loan had gained $1,845.00 after 36 months, what is the simple interest rate?

Answer by Boreal(15235)

I=prt
interest is 1845=7500*r*3 for 3 years
r=1845/3*7500=0.082 or 8.2%